R has a ten-year loan of $95685 at 6% interest. What payment will R make in the seventh year?

 


1
Expert's answer
2021-09-28T18:13:12-0400

"A=P(1+r\/100)^n"

"r=6\%"

first year, A"=" "95685\times 1.06=101426.1"

2nd year, "A=101426.1\times1.06=107511.666"

3rd, "A=107511.666\times 1.06=113962.366"

4th year, "A=113962.366\times 1.06=120800.1079"

5th year, "A=120800.1079\times =128048.1144"

6th year, "A=128048.1144\times 1.06=13573.0013"

7th year, "A=135731.0013\times 1.06=\$143874.8613"